OpenStudy (riley):

Find the marginal cost, marginal revenue, and marginal profit functions. Find all values of x for which the marginal profit is zero. C(x) = 5x R(x) = 8x − 0.001x^2

OpenStudy (zarkon):

find C'(x) R'(x) and P'(x)=R'(x)-C'(x) solve P'(x)=0 for x

OpenStudy (riley):

okay hmmm lemme see

OpenStudy (riley):

C'(x) = 5 R'(x) = 8-0.002x

OpenStudy (riley):

so then P'(x) = 3-0.002x right?

OpenStudy (zarkon):

yes

OpenStudy (riley):

okay so then I just plug in 0? O_O

OpenStudy (zarkon):

solve 3-0.002x=0

OpenStudy (riley):

oh oh oh, I set it equal to zero!! I gotcha now

OpenStudy (riley):

so that would come out to be 3 = 0.002x and then I divide 3 by 0.002, so x =

OpenStudy (zarkon):

yes

OpenStudy (riley):

1500 o.o

OpenStudy (zarkon):

1500

OpenStudy (riley):

yay xDDDDD ty ty ty Zarkon!!<3

OpenStudy (zarkon):

np